# Sparse Mamba: Reinforcing Controllability In Structural State Space Models

### v2

@@ Abstract
In this work, we introduce the concept of controllability and observability to the Mamba SSM’s architecture in our Sparse-Mamba (S-Mamba) for natural language processing (NLP) applications. The structured state space model (SSM) development in recent studies, such as Mamba and Mamba2, outperformed and solved the computational inefficiency of transformers and large language models at small to medium scale. The Mamba SSMs architecture drops the need for attention layers or multilayer perception blocks in transformers. However, current Mamba models lack reinforcement of controllability in state-space equations for computing the $A$, $B$, $C$, and $D$ matrices at each time step, leading to increased complexity and computational costs. In this paper, we demonstrate a reduction of parameters in comparison to the first published Mamba and Mamba2. We showcase an improvement in perplexity by 5% and a decrease in training time by 3% after reinforcing controllability and observability on the original Mamba architecture in our proposed S-Mamba. The controllable $n\times n$ state matrix $A$ is sparse and it has only $n$ free parameters. Our novel approach will ensure a controllable system which will be the gate key for Mamba3.
@@ 1 Introduction
In this work, we introduce a new family of sparse SSMs based on the fundamental control theory concepts of controllability and observability developed by Kalman in 1960s. In particular, we investigate how vanilla Mamba overlooked important concepts in control theory: controllability and observability. Therefore, we propose a family of Sparse Mamba (S-Mamba) networks where a modification on the architecture of vanilla Mamba can reinforce the system to be in the controller canonical form and in the observable canonical form [20]. Discussed in details in section [3.2,4], S-Mamba outperforms the original Mamba, reduces the number of parameters, and saves time in training. We start presenting our work by explaining the core structure of SSMs in Section [2]. Then, we will explain the building blocks, $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices in particular, of the vanilla Mamba and our S-Mamba in Section [3]. We evaluate our work in Section [4] and present the results in Tables [1,2,3].
@@ 3 Mamba
The structure of state space representations in $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices has an enormous impact on the SSM’s performance. Furthermore, the initialization of these matrices is as critical. Discussion in Section [2] revolved specifically around the building blocks of Mamba and around SSMs in general. Here, we present our novel method of initialization and calculation of $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ in our sparse Mamba S-Mamba. This presentation is done through showing these matrices’ structure in Mamba first, then ours afterwords. From this point on, mentioning Mamba will refer to vanilla Mamba version as Mamba [17] and S-Mamba will refer to the family of sparse mamba: Controlable Mamba SC-Mamba and Observable Mamba SO-Mamba.
@@ 3.2 Sparse Mamba Using Controllable and Observable Forms
In control theory, the controllable canonical form is a specific configuration of state-space representation where the state matrix $\mathbf{A}$, the input matrix $\mathbf{B}$, and the output matrix $\mathbf{C}$ have specific structured forms. The $n\times n$ matrix $\mathbf{A}$ is arranged in such a way that it makes the system controllable [24], and this form is particularly useful for state feedback control design discussed in section [3.2.1]. We will further implement and discuss the observable form [25] in section [3.2.2].
@@ 3.2.1 Controllability
The first part of our Sparce Mamba Family is Sparce Controllable Mamba (SC-Mamba). Here, the derivation of the controllable canonical form $CCF$ is closely related to the concept of reachability. A system is said to be reachable if it is possible to drive the state from any initial state to any final state within a finite time interval using an appropriate control input. In other words, The system is reachable if and only if the reachability matrix $R$ has full rank [20]. The CCF makes the system’s controllability properties explicit. This means that it is easier to analyze and design controllers for the system because the controllability matrix is in a specific, structured form. Furthermore, since the CCF provides a clear structure, it simplifies the design of state feedback controllers. The placement of poles and zeros becomes more manageable. Here, a linear time-invariant system represented by the transfer function 17. The state matrix $\mathbf{A}$ in controllable canonical form is structured as Eq.(18).
@@ 4 Experimental Results
Mamba 10.4618 99.2525 70.33 30.5017
SO-Mamba 10.0525 99.3704 72.2705 30.1241
SC-Mamba 9.8904 98.5427 69.6179 30.0191
@@
In Table [1], we present the improvement in perplexity in our family Sparse Mamba. Here, the sparse controllable Mamba (SC-Mamba) shows an improvement of 5% in comparison to the original vanilla Mamba model. Additionally, we mention the results of enforcing the observability matrix in SO-Mamba. Table [2], shows a reduction by 3% in training time. The training for the models on each of the datasets was done for $7$ epochs and the comparison was made on the last epoch.
@@ 5 Conclusion
In our paper, we introduce a family of Sparse Mamba (S-Mamba) that reinforces the controllability and observability on the original Mamba model. The controllable $n\times n$ state matrix $A$ is sparse and it has only $n$ free parameters. We showcase that our novel architecture has a better performance in terms of perplexity, less training time, and a reduction in the number of parameters in than vanilla Mamba. Our experiments prove a possibility to make any model that is based on state space representations sparse, including the diagonal structure in Mamba2, by enforcing Controllability and observability in $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices. This will conclude a less complex system for language modeling using SSMs.

### v3

# Sparse Mamba: Introducing Controllability, Observability, And Stability To Structural State Space Models

@@ Abstract
Structured state space models’ (SSMs) development in recent studies, such as Mamba and Mamba2, outperformed and solved the computational inefficiency of transformers and large language models at small to medium scale. In this work, we introduce the concept of controllability and observability to the original Mamba SSM’s architecture in our Sparse-Mamba (S-Mamba) for natural language processing (NLP) applications. Moreover, we reinforce stability on the $nxn$ $A$ matrix on Mmaba2. The Mamba SSMs architecture drops the need for attention layers or multilayer perception blocks in transformers. However, current Mamba models lack reinforcement of controllability in state-space equations for computing the $A$, $B$, $C$, and $D$ matrices at each time step, leading to increased complexity and computational costs. Furthermore, the $A$ matrix in Mamba2 is not always stable. We demonstrate a reduction of parameters compared to the first published Mamba and Mamba2. We showcase an improvement in perplexity by 5% and a decrease in training time by 3% after reinforcing controllability and observability on the original Mamba architecture in our proposed S-Mamba. We further enforce stability on the $A$ matrix in Mamba2 to improve the loss and perplexity of the model. The controllable and stable $n\times n$ state matrix $A$ is sparse, and it has only $n$ free parameters. Our novel approach will ensure controllable/observable and stable SSMs, which will be the gate key for Mamba3.
@@ 1 Introduction
In this work, we introduce a new family of sparse SSMs based on the fundamental control theory concepts of controllability and observability developed in [10]. In particular, we investigate how vanilla Mamba overlooked important concepts in control theory: controllability and observability. We further enforce the stability structure on the $A$ matrix in Mmaba2 [20]. Therefore, we propose a family of Sparse Mamba (S-Mamba) networks where a modification on the architecture of vanilla Mamba can reinforce the system to be in the controller canonical form and in the observable canonical form [21] on Mamba and the system to be stable in Mamba2. Discussed in details in section [3.3,4], S-Mamba outperforms the original Mamba, reduces the number of parameters, and saves time in training. We start presenting our work by explaining the core structure of SSMs in Section [2]. Then, we will explain the building blocks, $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices in particular, of the vanilla Mamba and our S-Mamba in Section [3]. We evaluate our work in Section [4] and present the results in Tables [1,2,3].
@@ 3 Mamba
The structure of state space representations in $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices has an enormous impact on the SSM’s performance. Furthermore, the initialization of these matrices is as critical. Discussion in Section [2] revolved specifically around the building blocks of Mamba and around SSMs in general. Here, we present our novel method of initialization and calculation of $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ in our sparse Mamba S-Mamba. This presentation is done through showing these matrices’ structure in Mamba first, then ours afterwords. From this point on, mentioning Mamba will refer to vanilla Mamba version as Mamba [17], Mamba2 will refer to the second version of Mamba [19] and S-Mamba will refer to the family of sparse mamba: Controlable Mamba (SC-Mamba), Observable Mamba (SO-Mamba), and Stable Mamba (ST-Mmaba2).
@@ 3.1 Mamba
3.2 Mamba 2
In a selective State Space Model (SSM), the state transition matrix $\mathbf{A}$ is time-dependent and has a shape of $\mathbf{A}\in\mathbb{R}^{(T,N,N)}$, where $T$ is the sequence length, and $N$ is the state dimension or size. Similar to Mmaba, to make computations efficient , Mamba2 restricts $\mathbf{A}$ to a diagonal structure, which reduces the shape to $(T,N)$ by storing only the diagonal elements of the $N\times N$ matrices. However, unlike the HiPPO matrix form, the $\mathbf{A}$ parameter in Mamba2 is further simplified to a scalar times the identity matrix, meaning all diagonal elements are the same value. In this form, $\mathbf{A}$ can be represented with shape $(T)$, treating $\mathbf{A}_{t}$ as a single scalar value, $a_{t}$ as shown in the matrix [17], where the elements $a_{t}$s in vector $\mathbf{A}$ is a uniformly distributed over a predefined range higher than 0.
$\mathbf{A}=\mathbf{-}\operatorname{diag}(a_{1},a_{2},a_{3},\dots,a_{t})$ (17)
The matrices $\mathbf{B}$ and $\mathbf{C}$ in a $Mamba2$ also vary with time, allowing for more flexibility in modeling temporal dependencies. The input matrix $\mathbf{B}$ has a shape of $\mathbf{B}\in\mathbb{R}^{(T,N)}$, while the output matrix $\mathbf{C}$ has the same shape, $\mathbf{C}\in\mathbb{R}^{(T,N)}$. These shapes imply that both $\mathbf{B}$ and $\mathbf{C}$ adapt for each time step in the sequence, giving the model fine-grained control over how the input $x_{t}$ influences the hidden state $h_{t}$ and how the state is mapped to the output $y_{t}$. This flexibility allows the SSM to selectively filter information, enabling better compression and state representation for time series modeling. Finaly, the $\mathbf{D}$ matrix is initialized as a vector of 1’s and set to be a learnable parameter same as $Mamba$.
@@ 3.3 Sparse Mamba Using Controllable and Observable Forms
In control theory, the controllable canonical form is a specific configuration of state-space representation where the state matrix $\mathbf{A}$, the input matrix $\mathbf{B}$, and the output matrix $\mathbf{C}$ have specific structured forms. The $n\times n$ matrix $\mathbf{A}$ is arranged in such a way that it makes the system controllable [25], and this form is particularly useful for state feedback control design discussed in section [3.3.1]. We will further implement and discuss the observable form [26] in section [3.3.2].
@@ 3.3.1 Controllability
The first part of our Sparce Mamba Family is Sparce Controllable Mamba (SC-Mamba). Here, the derivation of the controllable canonical form $CCF$ is closely related to the concept of reachability. A system is said to be reachable if it is possible to drive the state from any initial state to any final state within a finite time interval using an appropriate control input. In other words, The system is reachable if and only if the reachability matrix $R$ has full rank [21]. The CCF makes the system’s controllability properties explicit. This means that it is easier to analyze and design controllers for the system because the controllability matrix is in a specific, structured form. Furthermore, since the CCF provides a clear structure, it simplifies the design of state feedback controllers. The placement of poles and zeros becomes more manageable. Here, a linear time-invariant system represented by the transfer function 18. The state matrix $\mathbf{A}$ in controllable canonical form is structured as Eq.(19).
@@ 3.3.2 Observability
3.3.3 Stable Mamba2
In Mamba2 architecture [19], authors multiply the diagonal $A$ matrix in the forward process in Eq.[17]. The multiplication ensures each entry along the diagonal of $A$ is non-positive assuming all the enters are positive. However, the A matrix is stable only if all eigenvalues of $A$ are negative real numbers or have negative real parts to complex number eigenvalues [28]. In our Stable Mamba2 (ST-Mamba2), we assert stability by selecting only the positive entries in the $A$ matrix and convert them to $a_{i}=-1\times 10^{-5}$. In Eq.[25], each element is tested and modified conditionally: positive and zero values are set to a small negative number and only inherently negative values remain unchanged. This added conditionality directly controls the eigenvalue behavior, which reinforce stability.
$a_{i}=\begin{cases}a_{i}&\text{, if }a_{i}<0,\\ -1\times 10^{-5}&\text{, if }a_{i}\geq 0.\end{cases}$ (25)
In state-space models where system stability is critical for producing reliable and bounded outputs [20]. By enforcing stability at the matrix level, our implementation prevents divergence in state trajectories, especially in iterative or recursive processes where the system state could otherwise grow unbounded. This makes the model more robust and predictable under various initial conditions and parameter settings. Additionally, setting zero values to a small positive number avoids potential issues with singular matrices or undefined dynamics, while keeping eigenvalues in the stable region.
@@ 4 Experimental Results
Mamba 10.46 99.25 70.33 30.50
SO-Mamba 10.05 99.37 72.27 30.12
SC-Mamba 9.89 98.54 69.62 30.02
@@
Table 4: Perplexity Evaluation Table: Training results comparison between Mamba2 and our sparse stable Mamba2 (ST-Mamba2) based on perplexity matrix. Numbers in parentheses, (1M) and (100K), stand for the number of rows used in each of the datasets.
Model CodeParrot 1M OpenWebText 1M Cosmopidia 100K
Mamba2 7.87 96.65 30.61
ST-Mamba2 7.53 96.46 29.74
@@
In Table [1], we present the improvement in perplexity in our family Sparse Mamba. Here, the sparse controllable Mamba (SC-Mamba) shows an improvement of 5% in comparison to the original vanilla Mamba model. Additionally, we mention the results of enforcing the observability matrix in SO-Mamba. Table [2], shows a reduction by 3% in training time. The training for the models on each of the datasets was done for $7$ epochs and the comparison was made on the last epoch. We further compare the performance between Mmaba2 and our stable Mamba2 (ST-Mamba2) in Table [4]. We prove the improvement in terms of perplexity when enforcing the stability on the $A$ matrix in Mamba2 architecture.
@@ 5 Conclusion
In our paper, we introduce a family of Sparse Mamba (S-Mamba) that reinforces the controllability/observability on the original Mamba model and stability on Mamba2 model. The controllable/observable and the stable $n\times n$ state matrix $A$ is sparse and it has only $n$ free parameters. We showcase that our novel architecture has a better performance in terms of perplexity, less training time, and a reduction in the number of parameters in than vanilla Mamba. Our experiments prove a possibility to make any model that is based on state space representations sparse, including the diagonal structure in Mamba2, by enforcing Controllability and observability in $(\mathbf{A},\mathbf{B},\mathbf{C},\mathbf{D})$ matrices. This will conclude a less complex system for language modeling using SSMs.